OpenStudy (anonymous):

(-2x^3y^5)(-3x^4y)^3 simplify expression

OpenStudy (mysesshou):

\[(-2x^3y^5)(-3x^4y)^3\] = (-2x^3y^5)((-3)^3x^12y^3) =(-2)(-3)^3*x^15*y^8 =(-2)(-27)*x^15*y^8 =(54)*x^15*y^8 i think